# Fedor V. Fomin

### v1

@@ 24            $\mathcal{M}\leftarrow\mathcal{M}\cup\{(M,f)\}$;
28if $\mathcal{M}\neq\emptyset$ then
@@
If $H_{i}$ is not empty, a call to Lemma 2 is made in Algorithm 1. By 7, the call is valid, and therefore $O$ is either an independent set of size $k$ in $G[C_{i}]$, or the subgraph $M_{i}$ in the desired TM-embedding. Since the condition statement before Algorithm 1 explicitly verifies that $O$ is not an independent set of size $k$, in Algorithm 1 the latter case holds. ∎
@@
First, we argue that $M$ is a subgraph of $G$ by considering three types of edges in $M$. Edges with two endpoints in $S$ are not changed between $M_{W}$ and $M$, and therefore are present in $G$ since $M_{W}[S]$ is a subgraph of $G$ by CD4. Consider an edge between $s\in S$ and $v\in C_{i}$, $i\in[t]$, in $M$. By construction of $M$, there exists an edge between $s$ and $w\in W_{i}$ in $M_{W}$ with $\tau(w)=v$. By CD5, $N_{M_{W}}(w)\cap S\subseteq N_{G}(\tau(w))$, meaning that $s$ and $v=\tau(w)$ are adjacent in $G$. Finally, consider an edge of $M$ with both endpoints outside $S$. By construction, the edge is a part of $M_{i}$, for some $i\in[t]$, and is present in $G$ since $M_{i}$ is a subgraph of $G[C_{i}]$.
@@
Proof of Claim 10. Consider the list TM-embedding $(M^{*},f^{*})$ of $H$ in $G$ of maximum size. By 5, there exists a cut descriptor $(W,M_{W},\xi,f_{W})$ with $M_{W}[S]=M^{*}[S]$, and $M^{*}$ hits (touches/avoids) $C_{i}$ if and only if $M_{W}$ hits (touches/avoids) $W_{i}$ for each $i\in[t]$. Consider the TM-embedding $(M,f)$ constructed by the algorithm at Algorithm 1 with the cut descriptor set to $(W,M_{W},\xi,f_{W})$. Note that since the algorithm reached Algorithm 1, all calls to Lemma 2 resulted in a TM-model, and $M$ is successfully constructed. By construction, $M[S]=M_{W}[S]=M^{*}[S]$. Observe also that if $M_{W}$ hits $W_{i}$, then the TM-model $M_{i}$ returned by Lemma 2 spans $C_{i}$, meaning $|M\cap C_{i}|\geq|M^{*}\cap C_{i}|$. If $M^{*}$ touches $C_{i}$, then $|M^{*}\cap C_{i}|=|W_{i}|$, and $M$ has at least as many vertices in $C_{i}$. If $M^{*}$ avoids $C_{i}$, $|M^{*}\cap C_{i}|=0$. Therefore, in each $S$, $C_{1}$, $C_{2}$, …, $C_{t}$, $M$ has at least as many vertices as $M^{*}$, meaning $|M|\geq|M^{*}|$, proving the claim. ∎
@@
Proof of Claim 11. Assume a list TM-model of $H$ in $G$ exists. Then by 5, there exists also a cut descriptor. Consider the iteration of the algorithm with this cut descriptor. Since the algorithm reaches Algorithm 1, all calls to Lemma 2 resulted in a suitable TM-embedding being found, therefore the algorithm necessarily adds a TM-embedding $(M,f)$ to $\mathcal{M}$ at Algorithm 1. This contradicts the assumption that Algorithm 1 is reached. ∎
Finally, we estimate the running time. First note that all lines except Algorithm 1 and Algorithm 1 incur at most a polynomial factor. By Lemma 2, up to polynomial factors Algorithm 1 takes time $2^{(h+s+k)^{\mathcal{O}(k)}}$, for each invocation. This holds since $|V(H_{i})|+|E(H_{i})|\leq 3h+3s$, see the proof of 7. Moving to Algorithm 1, let $w=h+2s$ be the maximum size of $W$; there are at most $w^{t}$ choices for $\xi$, and at most $|V(H)|^{s+w}$ choices for $f_{W}$. We may assume $t\leq k$, since otherwise, an independent set of size $k$ can be found trivially by taking an arbitrary vertex from each $C_{i}$. We now argue that with a fixed $f_{W}$, there are at most $(s+w)^{2(s+w)}$ choices for $M_{W}$ that satisfy CD1. Indeed, terminals are fixed by $f_{W}$, therefore, it only remains to set the edges where at least one endpoint is a non-terminal. Since every non-terminal has degree two in $M_{W}$, there are at most $(s+w)^{2}$ choices for each non-terminal, and at most $(s+w)^{2(s+w)}$ choices in total. This part can also be upper-bounded by $(s+h)^{\mathcal{O}(s+h)}=2^{(s+h)^{1+o(1)}}$. The overall running time is therefore at most $2^{(s+h+k)^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}$. ∎
@@ 2.3 Proof of Theorem 1
1 $S\leftarrow\emptyset$; for $j\leftarrow 1$ to $\infty$ do
2       $C_{1}$, …, $C_{t}\leftarrow$ connected components of $G\setminus S$; if $t\geq k$ then
3             $I\leftarrow\{v_{1},\ldots,v_{k}\}$, where $v_{i}\in C_{i}$; return $I$.
@@ 2       $C_{1}$, …, $C_{t}\leftarrow$ connected components of $G\setminus S$; if $t\geq k$ then
5      $\kappa_{j}\leftarrow 3\cdot\max\{k+2,10\}\cdot(|V(H)+|E(H)|+|S|)$; if for each $i\in[t]$, $C_{i}$ is $\kappa_{j}$-connected then
@@
8             pick $i$ s.t. $C_{i}$ is not $\kappa_{j}$-connected; find minimum separator $S^{\prime}$ in $G[C_{i}]$; $S\leftarrow S\cup S^{\prime}$;
@@
3 Proof of Theorem 2: Path cover bellow Gallai-Milgram
In this section, we prove Theorem 2. Recall that in the Path Cover problem, we are given a graph $G$ and integer $p\geq 1$, and the task is to decide whether it is possible to cover $G$ with at most $p$ vertex-disjoint paths. Throughout this section, by $\operatorname{pc}(G)$, we denote the minimum number of vertex-disjoint paths required to cover $G$. A well-known theorem by Gallai and Milgram [GM60] guarantees that $\operatorname{pc}(G)\leq\alpha(G)$. When we combine this theorem with Theorem 1, we obtain an algorithm solving Path Cover in time time $2^{\alpha(G)^{\mathcal{O}(\alpha(G)^{2})}}\cdot|G|^{\mathcal{O}(1)}$. Our contribution in Theorem 2 makes a step further by establishing the tractability of Path Cover for parameterization $k=\alpha(G)-p$, that is “below” the bound of Gallai-Milgram’s theorem. We restate Theorem 2 here for the reader’s convenience.
See 2
We require two auxiliary results to prove Theorem 2. The first result is dealing with small vertex separators of $G$. It provides a useful tool that for a small vertex separator $S$ of $G$, reduces the number of connected components in $G-S$ and the value of $\operatorname{pc}(G)$ simultaneously.
Lemma 4.
Proof.
Claim 13.
Proof of Claim 13. Notice that if $\mathcal{P}$ has no path with an edge $uv$ with $u\in S$ and $v\in C_{i}$ for some $i\in[t]$ then the clique $C_{i}$ is covered by a degenerate path by the minimality of $\mathcal{P}$. Hence, if $C_{i}$ contains a vertex of a non-degenerate path then $\mathcal{P}$ has a path $P$ with an edge with one endpoint in $S$ and the other in $C_{i}$. Each vertex in $S$ is incident to at most two edges in the paths from $\mathcal{P}$. Thus, at most $2|S|$ edges are in the path of $\mathcal{P}$ with one endpoint in $S$ and the other in one of $C_{1},C_{2},\ldots,C_{t}$. Hence, at most $2|S|$ of the sets can contain vertices of non-degenerate paths. ∎ To construct $X$, for each vertex $v\in S$, we mark at most $2|S|$ cliques from $\{C_{i}\}_{i\in[t]}$ that contain neighbors of $v$. If there are at most $2|S|$ such cliques, then we mark all of them. Otherwise, we mark arbitrary $2|S|$ cliques containing neighbors of $v$. Then $X$ is defined as the set of indices of the marked cliques. It is straightforward that $|X|\leq 2|S|^{2}$ and $X$ can be constructed in polynomial time. We claim that there is a minimum path cover $\mathcal{P}$ of $G$ such that for each $i\in[t]\setminus X$, there is a path $P\in\mathcal{P}$ satisfying $V(P)=C_{i}$. In other words, each unmarked clique is covered by a degenerate path. Given a path cover $\mathcal{P}$ of $G$, we say that an edge $uv$ of a path $P\in\mathcal{P}$ is bad if $u\in S$ and $v$ is a vertex of an unmarked clique. We show that there is a minimum path cover $\mathcal{P}$ of $G$ that does not contain paths with bad edges. The proof is by contradiction. We choose $\mathcal{P}$ as a minimum path cover of $G$ that minimizes the total number of bad edges in the paths. Suppose that there is $P\in\mathcal{P}$ that contain a bad edge $uv$ with $u\in S$ and $v\in C_{i}$ for some $i\in[t]$. We have that $C_{i}$ is unmarked and has a neighbor of $u$. Then, by the definition of the marking procedure, there are at least $2|S|$ marked cliques with neighbors of $u$. Because $C_{i}$ contains a vertex of the non-degenerate path $P$, we obtain that there is a marked clique $C_{j}$ for some $j\in[t]$ with a neighbor $w$ of $u$ such that $C_{j}$ is covered by a degenerate path $P^{\prime}\in\mathcal{P}$ by 13. Because $C_{j}$ is a clique, we can assume without loss of generality that $w$ is an end-vertex of $P^{\prime}$. Deleting the edge $uv$ separates $P$ into two paths $Q$ and $Q^{\prime}$ having end-vertices $v$ and $u$, respectively. We use the fact that $uw\in E(G)$ and construct the path $\hat{Q}$ from $Q^{\prime}$ and $P^{\prime}$ by joining their end-vertices $u$ and $w$ by the edge $uw$. Notice that $Q$ and $\hat{Q}$ are vertex-disjoint and $V(Q)\cup V(\hat{Q})=V(P)\cup V(P^{\prime})$. This allows us to construct the path cover $\mathcal{P}^{\prime}$ by replacing $P$ and $P^{\prime}$ in $\mathcal{P}$ by $Q$ and $\hat{Q}$. Because the number of paths in $\mathcal{P}^{\prime}$ is the same as in $\mathcal{P}$, $\mathcal{P}^{\prime}$ is a minimum path cover. However, the number of bad edges in the path of $\mathcal{P}^{\prime}$ is less than the number of bad edges in the paths of $\mathcal{P}$, contradicting the choice of $\mathcal{P}$. This proves that there is a minimum path cover $\mathcal{P}$ of $G$ that does not contain paths with bad edges. To complete the proof, we remind that if a minimum path cover has no path with an edge $uv$ with $u\in S$ and $v\in C_{i}$ for some $i\in[t]$ then the clique $C_{i}$ is covered by a degenerate path. Thus, there is a minimum path cover $\mathcal{P}$ of $G$ such that for each $i\in[t]\setminus X$, there is a path $P\in\mathcal{P}$ satisfying $V(P)=C_{i}$. ∎
Our following auxiliary result is a “lighter version” of Theorem 2. It deals with the particular case of $\operatorname{pc}(G)<k^{C}$.
Lemma 5.
Proof.
We present an algorithm $\mathcal{A}$. The algorithm assumes that $G$ contains at least one edge as, otherwise, $pc(G)=n$ and the minimum-size path cover consists of trivial paths. Given $G,k$ and a path cover $\mathcal{P^{\prime}}$ of $G$ of size $p^{\prime}$, the algorithm $\mathcal{A}$ tries to find an optimal path cover of $G$. To achieve that, the algorithm iterates an integer $i$ from $1$ up to $p^{\prime}$ and an integer $j$ from 00 to $i-1$. On each iteration, the algorithm constructs a graph $H_{ij}$ defined as the disjoint union of $j$ copies of $K_{1}$ and $i-j$ copies of $K_{2}$, that is, $G$ contains $j$ isolated vertices and $i-j\geq 1$ edges forming a matching. Notice that $G$ admits a path cover with $i$ paths if and only if $H_{ij}$ has a spanning TM-model in $G$ for some $j\in\{0,\ldots,i-1\}$ because $G$ has at least one edge and, therefore, at least one path in a path cover of minimum size is nontrivial. The algorithm $\mathcal{A}$ thus applies the algorithm of Theorem 1 as a subroutine to $H_{ij},G$ and trivial list assignment $L\equiv V(G)$, with integer parameter set to $p^{\prime}+k=\mathcal{O}(k^{C})$. The subroutine works in
running time. If the outcome of the subroutine is an independent set $I$ in $G$ of size $p^{\prime}+k$, then the algorithm outputs $\mathcal{P}^{\prime}$ and $I$ as the solution. Since $|\mathcal{P}^{\prime}|=p^{\prime}$, $|I|\geq|\mathcal{P}^{\prime}|+k$ is satisfied, so this solution satisfies the lemma statement. If the subroutine decides that $H_{ij}$ does not have TM-model in $G$ or the maximum TM-model of $H_{ij}$ in $G$ is not spanning, then $G$ cannot be covered with $i$ paths and $\mathcal{A}$ continues to the next iteration. If the subroutine gives a spanning TM-model of $H_{ij}$ in $G$, then the minimum path cover size of $G$ equals $i$ and $\mathcal{A}$ transforms this TM-model into a path cover of $G$ with $i$ paths. Then $\mathcal{A}$ outputs this optimal path cover of $G$ and stops. Note that $\mathcal{A}$ necessary stops and produces a correct solution at some iteration, as it iterates $i$ over $\{1,2,\ldots,p^{\prime}\}$ while $1\leq\operatorname{pc}(G)\leq p^{\prime}$. The proof of the lemma is complete. ∎
We are ready to prove the main result of this section.
Proof of Theorem 2.
We describe an algorithm that starts from a trivial path cover of $G$ and then gradually transforms it. These transformations are indicated as reduction rules and are performed in polynomial time. We show that each such rule does not increase the number of paths in the path cover. Moreover, each reduction rule necessarily reduces either the total number of paths or the number of special (to be defined later) paths in the path cover. A series of consecutive applications of the reduction rules always lead the algorithm to cases when the algorithm can produce a solution satisfying theorem statement in $f(k)\cdot n^{\mathcal{O}(1)}$ time and stop. We indicate such cases and the corresponding behavior of the algorithm as solution subroutines. The algorithm runtime thus always ends with a solution subroutine run. In what follows, we present the reduction rules and solution subroutines. Each of the reduction rules reduces the number of paths in a path cover or reduces the number of some special paths. Solution subroutines output path covers and independent sets. The rules and subroutines are given in the order in which the algorithm checks their applicability. That is, the algorithm applies a reduction rule or solution subroutine only if no reduction rule or solution subroutine before it can be applied. When we reach the situation that the obtained path cover $\mathcal{P}$ and graph $G$ that are irreducible, we deduce that the instance has nice structural properties. In this case, the final solution subroutine calls the algorithms of Lemmas 4 and 5.
Reduction Rule 1.
The correctness of 1 is trivial, and it reduces the number of paths in $\mathcal{P}$ by one. Clearly, the exhaustive application of 1 to $\mathcal{P}$ yields $m\leq\alpha(G)$, as $\{s_{1},s_{2},\ldots,s_{p}\}$ induces an independent set in $G$. When 1 cannot be applied, then the only edges in $G$ allowed between the endpoints of paths in $\mathcal{P}$ are the edges of form $s_{i}t_{i}$. Special and usual paths. We call a path $P_{i}$ in the path cover special, if either $s_{i}=t_{i}$, or $P_{i}$ consists of only edge $s_{i}t_{i}$, or its endpoints, are connected by the edge $s_{i}t_{i}$. If a path $P_{i}$ is not special, we call it usual. We claim that sufficiently many usual paths give a solution to the problem.
Claim 14.
Proof of Claim 14. We show that $G$ has an independent set $I$ of size $m+k$. Initialize $I$ with $\{s_{1},s_{2},\ldots,s_{p}\}$. Since 1 is not applicable, $I$ is independent in $G$. Let $i_{1},i_{2},\ldots,i_{k}$ be $k$ distinct indices of usual paths. Since each $P_{i_{j}}$ is usual, $t_{i_{j}}\neq s_{i_{j}}$ and $s_{i_{j}}$ is not connected with $t_{i_{j}}$. Also, 1 is not applicable, so there is no edge between $t_{i_{j}}$ and any vertex in $I$, while $\{t_{i_{1}},t_{i_{2}},\ldots,t_{i_{k}}\}$ is also an independent set in $G$. Put $t_{i_{1}},t_{i_{2}},\ldots,t_{i_{k}}$ in $I$. We obtain an independent set of size $m+k$ in $G$, consequently $\alpha(G)-k\geq m$. ∎ 14 yields the corresponding solution subroutine for the algorithm.
Solution Subroutine 1.
If there are at least $k$ usual paths among $\mathcal{P}=P_{1},P_{2},\ldots,P_{m}$, output $\mathcal{P}$ and $I$ as in 14.
Reduction Rule 2.
  1. 1.
    Add edge $s_{i}t_{i}$ to $P_{i}$, so it becomes a cycle;
  2. 2.
  3. 3.
    Join $P_{i}$ and $P_{j}$ via $us_{j}$ (or $ut_{j}$).
Reduction Rule 3.
  1. 1.
    Add edge $s_{j}t_{j}$ to $P_{j}$, so it becomes a cycle;
  2. 2.
    Remove edge $u_{j}v_{j}$ from $P_{j}$, so it becomes an $u_{j}v_{j}$-path.
  3. 3.
    Apply 2 to $P_{i}$, $P_{j}$ and $u_{i}u_{j}$.
Claim 15.
Proof of Claim 15. Assume that the reduction rules are not applicable, but there are paths $P_{i},P_{j}\in\mathcal{P}$ with $i\neq j$ with an edge between $u_{i}\in V(P_{i})$ and $u_{j}\in V(P_{j})$. If both $P_{i}$ and $P_{j}$ consist of more than two vertices, then 3 is applicable. Thus, without loss of generality, $P_{j}$ consists of one or two vertices, so $V(P_{j})=\{s_{j},t_{j}\}$ and $u_{j}\in\{s_{j},t_{j}\}$. If $P_{i}$ consists of more than two vertices, then 2 is applicable, which is a contradiction. Then, the length of $P_{i}$ is less than three, implying that $u_{i}\in\{s_{i},t_{i}\}$. In this case, 1 is applicable, leading us to the final contradiction. ∎ The claim is the basis for the next solution subroutine, which the algorithm applies next. We explain how it works by proving the following claim.
Claim 16.
Proof of Claim 16. We construct an independent set $I$ consisting of more than $m+k$ vertices. Since 1 is not applicable, there are more than $m-k$ special paths. For each special path $P_{i}$ such that $V(P_{i})$ is not a clique, put two independent vertices of $V(P_{i})$ in $I$. For each other special path, put its arbitrary vertex into $I$. By 15, $I$ is an independent set in $G$. Since there are at least $2k$ non-cliques among vertex sets of special paths, $|I|>(m-k)+2k=m+k$. ∎ The solution subroutine based on 16 is given below. The algorithm applies it when none of the reduction rules and solution subroutines above are applicable.
Solution Subroutine 2.
If there are at least $2k$ special paths $P_{i}$ such that $V(P_{i})$ is not a clique in $G$, then output $\mathcal{P}$ and an $I$ as in 16.
Reduction Rule 4.
  1. 1.
  2. 2.
  3. 3.
    Join $P_{i}$ with the prefix part via $u_{i}x_{i}$;
  4. 4.
    Join $P_{j}$ with the suffix part via $u_{j}x_{j}$.
Claim 17.
Proof of Claim 17. The second part of the claim is trivial, since 4 produces three (or two, if $x_{i}x_{j}\in E(P_{\ell})$) paths by transforming three distinct paths $P_{i},P_{j},P_{\ell}$, so $|\mathcal{P}|$ cannot increase. To prove the first part, we show that the resulting $v_{i}s_{\ell}$-path and $v_{j}t_{\ell}$-path are usual. We denote these two paths by $Q_{i}$ and $Q_{j}$ correspondingly. We give the proof only for $Q_{i}$, since the proof for $Q_{j}$ is symmetric. We first show that $|V(Q_{i})|\geq 3$. Since $\{u_{i},v_{i},x_{i},s_{\ell}\}\subset V(Q_{i})$, it is enough to show that there are at least three distinct vertices among $u_{i},v_{i},x_{i},s_{\ell}$. The only two pairs of vertices that can coincide are $u_{i},v_{i}$ and $x_{i},s_{\ell}$, since they belong to disjoint paths $P_{i}$ and $P_{\ell}$ correspondingly. Hence, if $|V(Q_{i})|<3$, then $u_{i}=v_{i}$ and $x_{i}=s_{\ell}$ both hold. But $u_{i}=v_{i}$ implies that $P_{i}$ has length 00 and $s_{i}=t_{i}=u_{i}$. Then $u_{i}x_{i}=s_{i}s_{\ell}$, so $s_{i}s_{\ell}\in E(G)$, and 1 is applicable for $P_{i}$ and $P_{\ell}$. This contradiction proves $|V(Q_{i})|\geq 3$. It is left to show that there is no edge between the endpoints of $Q_{i}$, $v_{i}$ and $s_{\ell}$. Indeed, if $v_{i}s_{\ell}\in E(G)$, then we have an edge between a special path $P_{i}$ and an endpoint of $P_{\ell}$, then 2 is applicable. This contradiction proves the claim. ∎
Claim 18.
If none of the Reduction Rules 1-5 and Solution Subroutines 1,2 can be applied to $\mathcal{P}$, then there exists $S\subseteq V(G)$ with $|S|<k$ such that $V(P_{i})$ is a connected component in $G-S$ for at least $m-2k$ special paths $P_{i}$. Moreover, such a set $S$ could be constructed in polynomial time.
Proof of Claim 18. Since 14 is not applicable, there are less than $k$ usual paths in $\mathcal{P}$. We construct $S$ by taking one or zero vertices from each usual path, so $|S|<k$ holds automatically. We now explain the choice of the vertex to put in $S$ for each usual path. Consider a usual path $P_{i}\in\mathcal{P}$. We call a vertex $u\in V(P_{i})$ a connector if there is a special path $P_{j}$ and a vertex $v\in V(P_{j})$ with $uv\in E(G)$. If $P_{i}$ has exactly one connector, then we put this connector in $S$. Otherwise, i.e. when $P_{i}$ has zero or more than one connectors, we do not put any vertex of $P_{i}$ in $S$. We show that at least $m-2k$ special paths that are isolated in $G-S$. First note that there are at least $m-k$ special paths in $\mathcal{P}$, since the number of usual paths is bounded. Assume now that $P_{j}$ is a special path that is not isolated in $G-S$. Then there is an edge $uv$ with $u\in V(P_{i})$, $v\in V(P_{j})$, $i\neq j$ and $u\notin S$. By 15, $P_{i}$ can only be usual. Since $u$ is a connector of $P_{i}$ and $u\notin S$, $P_{i}$ has two or more distinct connectors. As 4 is not applicable, all these connectors can only connect $P_{i}$ with $P_{j}$. Hence, each usual path $P_{i}$ gives at most one non-isolated special path in $G-S$. Thus, at most $k$ special paths in $G-S$ are not isolated, implying that at least $(m-k)-k=m-2k$ special paths are isolated in $G-S$. ∎
Solution Subroutine 3.
Perform the following steps:
  1. 1.
    Obtain $S$ according to 18 and put $h:=\max\{0,m-4k\}$;
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
    Output $\mathcal{S}$ and (if obtained) $I$.
Claim 19.
If Reduction Rules 1-5 and Solution Subroutines 1,2 are not applicable, then 3 works in $2^{k^{\mathcal{O}(k^{4})}}\cdot|G|^{\mathcal{O}(1)}$ time and produces a feasible solution.
Proof of Claim 19. First note that at least $m-4k$ complete connected components are guaranteed in $G-S$ by 18 ($S$ separates $m-2k$ special paths) and inapplicability of 2 (at most $2k$ vertex sets of special paths are not cliques). We now move on to establish the upper bound on the running time. All steps, except for applying Lemma 5, are carried out in polynomial time. We have
By applying Lemma 5 to $G^{\prime},k^{\prime}$ and $\mathcal{P}^{\prime}$, we obtain the running time $2^{k^{\mathcal{O}(k^{4})}}\cdot|G|^{\mathcal{O}(1)}$. Set $\mathcal{S}$ is obtained from a path cover $\mathcal{S^{\prime}}$ of $G^{\prime}$ via adding paths covering $V(G)\setminus V(G^{\prime})$. Hence it is also a path cover. On the other hand, $I$ is (if obtained) an independent set in $G$ because $I\subseteq V(G-S)$ and $I$ is independent in $G-S$. Now we prove that either $\mathcal{S}$ is optimal or $|I|-|\mathcal{S}|\geq k$. First, let us consider the scenario where $I$ is not produced. This can occur only when the algorithm of Lemma 5 does not produce $I^{\prime}$. Then Lemma 5 guarantees that $\mathcal{S^{\prime}}$ is an optimal vertex cover of $G^{\prime}$ and $|\mathcal{S}^{\prime}|=\operatorname{pc}(G^{\prime})$. By Lemma 4, removal of $V(P_{j})$ for any $j\in[h]\setminus X$ decreases $\operatorname{pc}(G)$ by one, so $\operatorname{pc}(G^{\prime})=\operatorname{pc}(G)-(h-|X|).$ Since $|\mathcal{P}\setminus\mathcal{P}^{\prime}|=h-|X|$, we have $|\mathcal{S}|=\operatorname{pc}(G)$. It means that $\mathcal{S}$ is an optimal path cover of $G$.
Finally, let us consider the situation where $I$ is produced. In this case, the algorithm in Lemma 5 provides $\mathcal{S^{\prime}}$ and $I^{\prime}$ with $|I^{\prime}-|\mathcal{S}^{\prime}|=k^{\prime}=2k$. We notice that
The feasibility of $\mathcal{S}$ and $I$ is proved. ∎
This completes the proof of Theorem 2. ∎

### v2

@@ 33            $\mathcal{M}\leftarrow\mathcal{M}\cup\{(M,f)\}$;
37
38if $\mathcal{M}\neq\emptyset$ then
@@
If $H_{i}$ is not empty, a call to Lemma 4 is made in Algorithm 1. By 14, the call is valid, and therefore $O$ is either an independent set of size $k$ in $G[C_{i}]$, or the subgraph $M_{i}$ in the desired TM-embedding. Since the condition statement before Algorithm 1 explicitly verifies that $O$ is not an independent set of size $k$, in Algorithm 1 the latter case holds. ∎
@@
First, we argue that $M$ is a subgraph of $G$ by considering three types of edges in $M$. Edges with two endpoints in $S$ are not changed between $M_{W}$ and $M$, and therefore are present in $G$ since $M_{W}[S]$ is a subgraph of $G$ by CD4. Consider an edge between $s\in S$ and $v\in C_{i}$, $i\in[t]$, in $M$. By construction of $M$, there exists an edge between $s$ and $w\in T_{i}$ in $M_{W}$ with $\tau(w)=v$. By CD5, $N_{M_{W}}(w)\cap S\subseteq N_{G}(\tau(w))$, meaning that $s$ and $v=\tau(w)$ are adjacent in $G$. Finally, consider an edge of $M$ with both endpoints outside $S$. By construction, the edge is a part of $M_{i}$, for some $i\in[t]$, and is present in $G$ since $M_{i}$ is a subgraph of $G[C_{i}]$.
@@
Proof of Claim 17. Consider the list TM-embedding $(M^{*},f^{*})$ of $H$ in $G$ of maximum size. By 12, there exists a cut descriptor $(W,M_{W},\xi,f_{W})$ with $M_{W}[S]=M^{*}[S]$, and $M^{*}$ hits (touches/avoids) $C_{i}$ if and only if $M_{W}$ hits (touches/avoids) $T_{i}$ for each $i\in[t]$. Consider the TM-embedding $(M,f)$ constructed by the algorithm at Algorithm 1 with the cut descriptor set to $(W,M_{W},\xi,f_{W})$. Note that since the algorithm reached Algorithm 1, all calls to Lemma 4 resulted in a TM-model, and $M$ is successfully constructed. By construction, $M[S]=M_{W}[S]=M^{*}[S]$. Observe also that if $M_{W}$ hits $T_{i}$, then the TM-model $M_{i}$ returned by Lemma 4 spans $C_{i}$, meaning $|M\cap C_{i}|\geq|M^{*}\cap C_{i}|$. If $M^{*}$ touches $C_{i}$, then $|M^{*}\cap C_{i}|=|T_{i}|$, and $M$ has at least as many vertices in $C_{i}$. If $M^{*}$ avoids $C_{i}$, $|M^{*}\cap C_{i}|=0$. Therefore, in each $S$, $C_{1}$, $C_{2}$, …, $C_{t}$, $M$ has at least as many vertices as $M^{*}$, meaning $|M|\geq|M^{*}|$, proving the claim. ∎
@@
Proof of Claim 18. Assume a list TM-model of $H$ in $G$ exists. Then by 12, there exists also a cut descriptor. Consider the iteration of the algorithm with this cut descriptor. Since the algorithm reaches Algorithm 1, all calls to Lemma 4 resulted in a suitable TM-embedding being found, therefore the algorithm necessarily adds a TM-embedding $(M,f)$ to $\mathcal{M}$ at Algorithm 1. This contradicts the assumption that Algorithm 1 is reached. ∎
Finally, we estimate the running time. First note that all lines except Algorithm 1 and Algorithm 1 incur at most a polynomial factor. By Lemma 4, up to polynomial factors Algorithm 1 takes time $2^{(h+s+k)^{\mathcal{O}(k)}}$, for each invocation. This holds since $|V(H_{i})|+|E(H_{i})|\leq 3h+3s$, see the proof of 14. Moving to Algorithm 1, let $w=h+2s$ be the maximum size of $W$; there are at most $w^{t}$ choices for $\xi$, and at most $|V(H)|^{s+w}$ choices for $f_{W}$. We may assume $t\leq k$, since otherwise, an independent set of size $k$ can be found trivially by taking an arbitrary vertex from each $C_{i}$. We now argue that with a fixed $f_{W}$, there are at most $(s+w)^{2(s+w)}$ choices for $M_{W}$ that satisfy CD1. Indeed, terminals are fixed by $f_{W}$, therefore, it only remains to set the edges where at least one endpoint is a non-terminal. Since every non-terminal has degree two in $M_{W}$, there are at most $(s+w)^{2}$ choices for each non-terminal, and at most $(s+w)^{2(s+w)}$ choices in total. This part can also be upper-bounded by $(s+h)^{\mathcal{O}(s+h)}=2^{(s+h)^{1+o(1)}}$. The overall running time is therefore at most $2^{(s+h+k)^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}$. ∎
@@ 4.3 Proof of Theorem 2
1
[…]
3for $j\leftarrow 1$ to $\infty$ do
4       $C_{1}$, …, $C_{t}\leftarrow$ connected components of $G\setminus S$;
5      if $t\geq k$ then
6             $I\leftarrow\{v_{1},\ldots,v_{k}\}$, where $v_{i}\in C_{i}$;
7            return $I$.
@@ 6             $I\leftarrow\{v_{1},\ldots,v_{k}\}$, where $v_{i}\in C_{i}$;
9
10      $\kappa_{j}\leftarrow 3\cdot\max\{k+2,10\}\cdot(|V(H)+|E(H)|+|S|)$;
11      if for each $i\in[t]$, $C_{i}$ is $\kappa_{j}$-connected then
@@
14             pick $i$ s.t. $C_{i}$ is not $\kappa_{j}$-connected;
15            find minimum separator $S^{\prime}$ in $G[C_{i}]$;
16            $S\leftarrow S\cup S^{\prime}$;